OpenStudy (anonymous):

If m, p and mp are the three real roots from x³ + mx² + mpx + p = 0 what's the sum of the roots? a)1 b)-1 c)2 d)3 I've tried to build a system of three equations and I got p=0 but the exercise says that none of the roots can be 0. What should I do? Pls help me.

OpenStudy (bahrom7893):

tricky lol.. thinking..

OpenStudy (anonymous):

maybe @ash2326 knows this one

OpenStudy (bahrom7893):

factor out something maybe.. x³ + mx² + mpx + p = 0 x²(x + m) + p(x + m) = 0

OpenStudy (bahrom7893):

i think i got this.. maybe lol (x+m)(x^2+p) = 0

OpenStudy (bahrom7893):

x = -m x^2 = -p, so p must be negative

OpenStudy (bahrom7893):

substitute x = -m into the equation...let's see what we get..

OpenStudy (bahrom7893):

x³ + mx² + mpx + p = 0 (-m)^3+m*(-m)^2 + m*p*(-m) + p = 0 -m^2*p + p = 0 p(1-m^2)=0

OpenStudy (ash2326):

Melinda what's the sum of the roots if the roots are x, y and x \[x^3+ax^2+bx+c=0\] x+y+z=?

OpenStudy (anonymous):

x³ + mx² + mpx + p = 0 x²(x + m) + p(x + m) = 0 => this doesnt look right

OpenStudy (anonymous):

m = - (m + mp + p) @ash2326

OpenStudy (bahrom7893):

yes it is melinda, pull out x^2 from the first two and then pull out a p from the last two

OpenStudy (bahrom7893):

m+p+m*p=0 p(1-m^2)=0 so either p = 0, or 1-m^2 = 0, but none of the solutions can be 0, so m^2 = 1 is one of the solutions, or, m can be either -1 or 1

OpenStudy (bahrom7893):

x^2 = -p, x = -m, so: (-m)^2 = -p

OpenStudy (bahrom7893):

m^2 = -p in either case, if u have 1 or minus one, m^2 is still 1, so -p = 1, or p=-1

OpenStudy (bahrom7893):

So p =-1, m = +/- 1

OpenStudy (bahrom7893):

m+p+mp = If m is -1: -1-1+(-1)(-1) = -2 + 1 = -1

OpenStudy (bahrom7893):

If m is 1: 1+(-1)+(1)(-1) = 1 - 2 = -1

OpenStudy (bahrom7893):

The sum is -1, elegant, eh? :D

OpenStudy (ash2326):

Yeah I'm also getting -1 , did you understand @MelindaR?

OpenStudy (anonymous):

you guys are awesome but actually i dont get it

OpenStudy (bahrom7893):

Okay Melinda I'll try to go step by step.

OpenStudy (ash2326):

I'm here, if you have any doubt, just ask for me

OpenStudy (anonymous):

Bah, I would factor x³ + mx² + mpx + p = 0 as x²(x + m) + p(mx + 1)=0

OpenStudy (bahrom7893):

Ohh shoot u're right, hold on then

OpenStudy (ash2326):

Melinda let me guide you

OpenStudy (anonymous):

the options are tricky, bah. xD ok ash

OpenStudy (ash2326):

We have the roots as m, p and mp so \[m+p+mp=-m\] \[mp+m^2p+mp^2=mp\] and \[m\times p\times mp=-p\] from third equation we get \[ m^2p^2+p=0 => p(m^2p+1)=0\] p can't be zero so \[m^2p=-1\] now in second equation we have \[mp+m^2p+mp^2=mp\] so we get \[m^2p+mp^2=0\] we have \[m^2p=-1\] so we get \[mp^2=-1\] so \[m^2p=-mp^2=> m^2p+mp^2=0\] we get \[ mp(p+m)=0\] mp can't be zero so \[ m=-p\] now use \[m+p+mp=-m\] m=-p sp \[m-m-m^2=-m\] we get \[ m^2-m=0\] \[m(m-1)=0\] m can't be zero so m=1 and p=-1 now \[mp=1\times -1\] \[mp=-1\] now sum of roots m+p+mp=1-1-1=-1

OpenStudy (bahrom7893):

ash howd u get the first three equations?

OpenStudy (anonymous):

its the theory of equations bah, check out this http://openstudy.com/study#/updates/4f4503f9e4b065f388dd4c1f

myininaya (myininaya):

\[x^3+mx^2+mpx+p=(x-m)(x-p)(x-m p)\] \[..................=(x^2-px-mx+m p)(x-m p)\] \[..................=(x^2+(-p-m)x+m p)(x-m p)\] \[..................=x^3+(-p-m)x^2+m p x-m p x^2-(-p-m)x m p -m^2p^2\] \[.................=x^3+(-p-m-m p)x^2+(m p +p^2+m)x -m^2p^2\] \[m=-p-m-m p ; m p=m p+p^2+m ; p=-m^2p^2\]

myininaya (myininaya):

If I didn't make a mistake...

OpenStudy (ash2326):

Thanks @MelindaR

OpenStudy (bahrom7893):

Ohhh, I see lol i didn't know that.. yay for another random fun fact lol

OpenStudy (bahrom7893):

all right myin's attacking the question.. so ima go back to probability

OpenStudy (anonymous):

Thanks a lot, Bah!

OpenStudy (ash2326):

Melinda did you understand?

OpenStudy (bahrom7893):

lol i didnt rly help haha, but u're welcome

OpenStudy (anonymous):

I'm trying ash, wait a sec pls

OpenStudy (anonymous):

I did all of that as well, ash... I'm trying to see where did i go wrong

OpenStudy (ash2326):

Take your time:)

myininaya (myininaya):

Yah... Ash I really like your answer. I can't top that icecream with whoop cream. Great Job! :)

myininaya (myininaya):

But at least I showed how to derive those equations you got :)

myininaya (myininaya):

So in a way I put sprinkles on it. :)

OpenStudy (ash2326):

Thanks @myininaya :D for extolling me. Yeah, you're the best, you made the ice cream heavenly:D

OpenStudy (anonymous):

@ash2326 I dont get this. so we get m2p+mp2=0 we have m2p=−1 so we get mp2=−1 if we have m²p = -1, then how come that m²p + mp² = 0 isn't - 1 + mp² = 0 and consequently mp² = 1?

OpenStudy (ash2326):

sorry I made a typo , you're right \[m^2p=-1 , mp^2=1\] and \[m^2p+mp^2=0\]

OpenStudy (anonymous):

@ash2326 Also, where did you get m² ? I only get that -m=p now use m+p+mp=−m m−m−m2=−m

OpenStudy (ash2326):

We have m+p+mp=-m and we got m=-p pr p=-m so \[ m+(-m)+m(-m)=-m\] we get \[m-m-m^2=-m\] or \[-m^2=-m=>-m^2+m=0=>m^2-m=0\]

OpenStudy (anonymous):

Oh, I can understand everything now. Thanks a lot @ash2326 and @myininaya !

OpenStudy (ash2326):

Welcome @MelindaR glad to help:D :D

myininaya (myininaya):

:)